OpenStudy (mackenzie2013):

The perimeter of a quadrilateral is 38 yards. If three of the sides measure 5.8 yards, 7 yards, and 11.2 yards, what is the length of the fourth side? 24 yd 62 yd 20.3 yd 14 yd

OpenStudy (mackenzie2013):

Wouldnt you add all of those up and subtract from 38?

OpenStudy (anonymous):

Okay since it is a quadrilateral it has 4 sides. Since you are given 3 sides and the perimeter you must find x. 38=5.8+7+11.2 sorry I have to go but yes you do that

OpenStudy (mackenzie2013):

Thank you!

OpenStudy (anonymous):

do 38 minus the the sum of the 3 sides. that should give you the unknown length

OpenStudy (anonymous):

38-(5.8+11.2+7)= 38-24= 14 D sorry I took a while I was taking my test=)

OpenStudy (anonymous):

so we r done. do u have more questions im happy to help:)